A soccer ball is kicked in the air the height h(t) of the ball in metres according to time in seconds is expressed as a function h(t) = -2t^2+8t. For what interval of t does the height of the ball increase??
Answer by stanbon(75887)

A soccer ball is kicked in the air the height h(t) of the ball in metres according to time in seconds is expressed as a function h(t) = -2t^2+8t. For what interval of t does the height of the ball increase??
---------
Vertex occurs at t = -b/(2a) = -8/(2*-2) = 2 seconds
------
Increase iinterval: (0,2)
-----------------------------
Cheers,
Stan H.